OpenStudy (anonymous):

List the solutions: A) sin x -1 = cosx B) sin(2x)-1=tanx

OpenStudy (anonymous):

I think the solution to A) is x=pi/2, x=pi

OpenStudy (perl):

a) sin(x) - 1 = cos(x), did I copy correctly?

OpenStudy (perl):

is that the question

OpenStudy (anonymous):

I'm not sure. It just says sin x - 1= coax, but I think you did

OpenStudy (perl):

how about squaring both sides

OpenStudy (anonymous):

could you show me? Are my answers to A) wrong?

OpenStudy (anonymous):

@perl

OpenStudy (perl):

(sin x -1 )^2 = (cos x) ^2 square both sides we get sin^2 (x) - 2*sin(x) + 1 = cos^2 (x)

OpenStudy (perl):

ok so far?

OpenStudy (anonymous):

I believe so

OpenStudy (perl):

square both sides gives us (sin x - 1 )(sin x - 1) = cos x * cos x then i foiled the left side

OpenStudy (perl):

so we have sin^2 (x) - 2*sin(x) + 1 = cos^2 (x) now on the right side, cos^2(x) , we can substitute 1 - sin^2(x) for it

OpenStudy (perl):

Here are the steps again: sin x - 1 = cos x (sin x - 1 )^2 = (cos x)^2 sin^2 (x) - 2*sin(x) + 1 = cos^2 (x) sin^2 (x) - 2*sin(x) + 1 = 1 - sin ^2 (x)

OpenStudy (anonymous):

Ok I get that so far

OpenStudy (perl):

we can substitute 1 - sin^2(x) for cos^2(x) since they are equal. do you see why they are equal? because sin^2 (x) + cos^2(x) = 1 , and subtract sin^2(x) from both sides

OpenStudy (perl):

cos^2(x) = 1 - sin^2(x) now is a new identity

OpenStudy (anonymous):

ok, so whats the next step @perl

OpenStudy (anonymous):

Ok I'm following so far @perl now we need to solve for x i think

OpenStudy (perl):

sin x - 1 = cos x (sin x - 1)^2 = (cos x)^2 sin^2 (x) - 2*sin(x) + 1 = cos^2 (x) sin^2 (x) - 2*sin(x) + 1 = 1 - sin ^2 (x) 2sin^2(x) - 2 sin(x) + 1 = 1 2 sin^2(x) - 2sin(x) = 0 2sin(x)*( sin (x) - 1) = 0 using zero product property 2 sin (x) = 0 OR sin(x) - 1 = 0 sin(x) =0 x = pi*n sin(x) -1 = 0 sin(x) = 1 x = pi/2 + 2pi*n

OpenStudy (anonymous):

oh so I was right lol?

OpenStudy (perl):

what is the domain , between 0 and 2pi ?

OpenStudy (anonymous):

yes

OpenStudy (perl):

actually the general solution is x = pi + 2pi*n, pi/2 + 2pi*n apparently it is not x = pi*n, since if n = 0 , x= pi*0 = 0 is a false solution

OpenStudy (anonymous):

oh ok so i;ll use the general solution answer. what about b)? I have to leave soon so please help like you just did @perl

OpenStudy (perl):

one second

OpenStudy (perl):

this one is tricky

OpenStudy (anonymous):

Yes I know. This is the one I really needed help on

OpenStudy (perl):

this one you have to use a calculator

OpenStudy (perl):

graph the two functions separately and estimate where they intersect

OpenStudy (perl):

use the zero option

OpenStudy (anonymous):

Could you give me the solution and show me how you got there? I really have no clue

OpenStudy (anonymous):

walk me through the steps would be great if you can. It's just I have to leave soon

OpenStudy (perl):

graph y = sin(2x)-1 y= tan(x)

OpenStudy (perl):

graph them separately,

OpenStudy (anonymous):

I have the graph, but it doesn't show me the solutions. could you tell me?

OpenStudy (perl):

sure

OpenStudy (perl):

do the directions say what the domain is ?

OpenStudy (anonymous):

should i include the 2pi*n on both parts? and yes 0,2pi

OpenStudy (perl):

if the problem says between 0 and 2pi, then don't include 2pi*n

OpenStudy (perl):

but you need to calculate 2pi ~ 6.28 so you want solutions between 0 and 6.28

OpenStudy (anonymous):

wait I'm confused now. I thought the answers were: A) x=pi/2, pi B)x= -0.59, x= 1.68

OpenStudy (anonymous):

the domain is [0,2pi] for both

OpenStudy (perl):

for a) x = pi/2, pi b) x = 2.0687813 , 5.210374

OpenStudy (anonymous):

oh ok. thanks for helping me here. Hopefully your correct :)

OpenStudy (perl):

does it say how much accuracy?

OpenStudy (anonymous):

no, just find the solutions

OpenStudy (perl):

I checked maple, by the way

OpenStudy (anonymous):

wait, again. what other the final solutions for B) . I see so many different answers

OpenStudy (perl):

oh shoot, that was a mistake again

OpenStudy (anonymous):

@perl so what is the final answer lol?

OpenStudy (anonymous):

\[\sin x-\cos x=1,squaring~ both~ sides\] \[\sin ^2x+\cos ^2x-2\sin x \cos x=1\] 1-sin 2x=1, sin 2x=0\[\sin 2x=0=\sin n \pi ,2x=n \pi,x=\frac{ n \pi }{ 2 }\] where n is an integer.

OpenStudy (anonymous):

why isn't this showing up for me?

OpenStudy (anonymous):

oh and @surjithayer I already know the answer to the first one. I need the 2nd

OpenStudy (perl):

sorry the program crashed

OpenStudy (anonymous):

@perl that;s ok. did you figure it out?

OpenStudy (perl):

a) x = pi/2, pi b) x = 2.0687813 , 5.210374

OpenStudy (queelius):

There's no exact solution to the second answer -- it'll be an irrational number. An approximate solution, of course, is possible. However, you seem to be confused by the notion that there are an infinite number of solutions, but the only solutions we are interested in is when they fall within some domain, namely, in this case, [0, 2pi].

OpenStudy (queelius):

Well, an irrational number that isn't a multiple of pi, and such, that is. :)

OpenStudy (anonymous):

Ok thanks guys. I believe @perl is correct for B) :)

OpenStudy (queelius):

Perl's answer is correct.

OpenStudy (perl):

here, I put it into wolfram http://www.wolframalpha.com/input/?i=solve+sin%282x%29-1%3D+tan%28x%29+%2C+0%3Cx%3C2pi

OpenStudy (queelius):

Just go to WolframAlpha and type: sin(2x)-1=tan(x).

OpenStudy (queelius):

Yup.

OpenStudy (queelius):

And look for solutions within the desired interval, e.g., sin(2x)-1=tan(x), from x=0 to x=2pi

OpenStudy (perl):

type: solve sin(2x)-1= tan(x) , 0<x<2pi

OpenStudy (perl):

queelius, wolf gives me an ugly answer if i dont specify x is between 0 and 2pi

OpenStudy (perl):

queelius, compare to this http://www.wolframalpha.com/input/?i=solve+sin%282x%29-1%3D+tan%28x%29+

OpenStudy (perl):

wolfram general solution is : x = 2 tan^(-1)(root of x^6+2 x^5+x^4-12 x^3-x^2+2 x-1 near x = -0.594554)+2pi *n x= 2 tan^(-1)(root of x^6+2 x^5+x^4-12 x^3-x^2+2 x-1 near x = 1.68193)+2pi*n

OpenStudy (queelius):

Perl, lol, that's when you ask for an exact answer. It gives you a nice formula dependent upon n, n any integer, but as you indicated, you need to choose an n for which the result falls within the desired interval. You can also just hover over the intersection points on the plot to get an approximate answer.

OpenStudy (perl):

thats a clue that there is no simple solution , in terms of pi or such

OpenStudy (perl):

what is interesting in part a) , when we squared both sides of the equation we introduced an extraneous solution

OpenStudy (anonymous):

ok guys lol is perls answer correct?

OpenStudy (perl):

yes wolfram concurs with me, and i did it using a TI 84, graphed and intersected

OpenStudy (anonymous):

ok thanks

OpenStudy (queelius):

Squaring both sides does allow for this: (-1)^2=(1)^2,

OpenStudy (queelius):

So you have to filter out extraneous solutions afterwards.

OpenStudy (queelius):

Btw, perl, earlier when I said "you seem to be confused by the notion that there are an infinute number..." I wasn't directing that to you, but to washcaps.

OpenStudy (queelius):

I just realized you may have thought I was referring to you. Random aside. :)